# Variational Continual Test-Time Adaptation

### v1

@@ 5.10 Hyperparameter analysis
Table 10: Different number of augmentation.
@@
6 Conclusion
To tackle the prior drift problem, in this paper, we proposed a variational Bayesian inference approach, termed VCoTTA, to estimate uncertainties in CTTA. At the source stage, we first transformed an off-the-shelf pre-trained deterministic model into a Bayesian Neural Network (BNN) using a variational warm-up strategy, thereby injecting uncertainty into the source model. At the test time, we implemented a mean-teacher update strategy, where the student model is updated via variational inference, while the teacher model is refined by the exponential moving average. Specifically, to update the student model, we proposed a novel approach that utilizes a mixture of priors from both the source and teacher models. Consequently, the evidence lower bound can be formulated as the cross-entropy between the student and teacher models, combined with the Kullback-Leibler (KL) divergence of the prior mixture. We demonstrated the effectiveness of the proposed method on three datasets, and the results show that the proposed method can mitigate the issue of prior drift within the CTTA framework.
Appendix A CTTA approximation by Bayesian inference (BI)
The goal of CTTA is to learn a posterior distribution $p(\boldsymbol{\theta}|\mathcal{U}_{1:T}\cup\mathcal{D}_{0})$ from a source dataset $\mathcal{D}_{0}$, and a sequence of unlabeled test data from $\mathcal{U}_{1}$ to $\mathcal{U}_{T}$. Following Zhou & Levine (2021), assuming we have multiple input-generating distributions that the source dataset $\mathcal{D}_{0}$ is drawn from a distribution $\phi$, and $\tilde{\phi}_{t}$ specifies the shifted of the $t$-th unlabeled test dataset which we aim to adapt to. Let the parameters of the model be $\boldsymbol{\theta}$,then following the semi-supervised learning framework (Grandvalet & Bengio, 2004), we incorporate all input-generating distributions into the belief over the model parameters $\boldsymbol{\theta}$ as follows
$p(\boldsymbol{\theta}|\phi,\tilde{\phi}_{1},\cdots,\tilde{\phi}_{T})$ $\propto p(\boldsymbol{\theta})\exp\left(-\lambda_{0}H_{\boldsymbol{\theta},\phi}(Y|\boldsymbol{X})\right)\prod_{t=1}^{T}\exp\left(-\lambda_{t}H_{\boldsymbol{\theta},\tilde{\phi}_{t}}(Y|\boldsymbol{X})\right),$ (20)
where the inputs $X$ are sampled i.i.d. from a generative model with parameters $\phi$, while the corresponding labels $Y$ are sampled from a conditional distribution $p(Y|X,\boldsymbol{\theta})$, which is parameterized by the model parameters $\boldsymbol{\theta}$. $p(\boldsymbol{\theta})$ is a prior distribution over $\boldsymbol{\theta}$. $\{\lambda_{0},\lambda_{1},\cdots,\lambda_{T}\}$ are the factors for approximation weighting. Generally, the entropy term $H_{\boldsymbol{\theta},\phi}(Y|\boldsymbol{X})$ represents the cross entropy of the supervised learning, and the entropy term $H_{\boldsymbol{\theta},\tilde{\phi}_{t}}(Y|\boldsymbol{X})$ for $t>0$ denotes the Shannon entropy of the unsupervised learning.
Following Zhou & Levine (2021), we can empirically use a point estimation to get a plug-in Bayesian approach to approximate the above formula:
$p(\boldsymbol{\theta}|\mathcal{U}_{1:T}\cup\mathcal{D}_{0})\propto p(\boldsymbol{\theta})\prod_{\forall x,y\in\mathcal{D}_{0}}p(y|x,\boldsymbol{\theta})\exp\left(-\frac{\lambda_{0}}{|\mathcal{D}_{0}|}\sum_{\forall x\in\mathcal{D}_{0}}{H}(Y|{x},\boldsymbol{\theta})\right)\prod_{t=1}^{T}\exp\left(-\frac{\lambda_{t}}{|\mathcal{U}_{t}|}\sum_{\forall x\in\mathcal{U}_{t}}{H}(Y|{x},\boldsymbol{\theta})\right).$ (21)
To make the formula feasible to CTTA, that is, no source data is available at the test time, we set $\lambda_{0}=0$. And the source knowledge can be represented by $p(\boldsymbol{\theta}|\mathcal{D}_{0})\propto p(\boldsymbol{\theta})\prod_{\forall x,y\in\mathcal{D}_{0}}p(y|x,\boldsymbol{\theta})$. Thus, for the $t$-th test domain, the Bayesian inference in CTTA can be represented as follows:
$p(\boldsymbol{\theta}|\mathcal{U}_{1:t}\cup\mathcal{D}_{0})$ $\propto p(\boldsymbol{\theta}|\mathcal{D}_{0})\prod_{i=1}^{t}\exp\left(-\frac{\lambda_{i}}{|\mathcal{U}_{i}|}\sum_{\forall x\in\mathcal{U}_{i}}{H}(Y|{x},\boldsymbol{\theta})\right)$ (22)
$p(\boldsymbol{\theta}|\mathcal{U}_{1:t}\cup\mathcal{D}_{0})\propto p(\boldsymbol{\theta}|\mathcal{U}_{1:t-1}\cup\mathcal{D}_{0})e^{-\lambda H(\mathcal{U}_{t}|\boldsymbol{\theta})}=p_{t}(\boldsymbol{\theta})e^{-\lambda H(\mathcal{U}_{t}|\boldsymbol{\theta})},$ (23)
Appendix B ELBO of the VI in CTTA
We built VI for CTTA in Sec. 3, where we initialize a variational distribution $q(\boldsymbol{\theta})$ to approximate the real posterior. For the test domain $t$, we optimize the variational distribution as follows:
where $\mathbb{Q}$ is the distribution searching space, and $p_{t}(\boldsymbol{\theta})$ is the current prior.
Following the definition of KL divergence and the standard derivation of the Evidence Lower BOund (ELBO) is as the following formulas. Specifically, the KL divergence is expanded as
$\mathrm{KL}\left[q(\boldsymbol{\theta})\parallel\cfrac{1}{Z_{t}}p_{t}(\boldsymbol{\theta})e^{-\lambda H(\mathcal{U}_{t}|\boldsymbol{\theta})}\right]$ $=-\int_{\boldsymbol{\theta}}q(\boldsymbol{\theta})\log\cfrac{\cfrac{1}{Z_{t}}p_{t}(\boldsymbol{\theta})e^{-\lambda H(\mathcal{U}_{t}|\boldsymbol{\theta})}}{q(\boldsymbol{\theta})}d\boldsymbol{\theta}$ (25)
@@ Appendix B ELBO of the VI in CTTA
where the first constant term can be reduced in the optimization. Thus, we can optimize the variational distribution via the ELBO:
$q_{t}(\boldsymbol{\theta})$ $=\arg\min_{q\in\mathbb{Q}}\mathrm{KL}\left[q(\boldsymbol{\theta})\parallel\frac{1}{Z_{t}}p_{t}(\boldsymbol{\theta})e^{-{\lambda}{H}(\mathcal{U}_{t}|\boldsymbol{\theta})}\right]$ (26)
$=\arg\max_{q\in\mathbb{Q}}-\lambda\mathbb{E}_{\boldsymbol{\theta}\sim q(\boldsymbol{\theta})}H(\mathcal{U}_{t}|\boldsymbol{\theta})-\mathrm{KL}\left(q(\boldsymbol{\theta})\parallel p_{t}(\boldsymbol{\theta})\right)$
$=\arg\max_{q\in\mathbb{Q}}\text{ELBO}.$
Optimization soving: In our case, the former entropy term can be more effectively replaced by the cross entropy or symmetric cross entropy (SCE) between the student model and the teacher model in a mean-teacher architecture (see Sec. 4.1). For the latter KL term, we can substitute a variational approximation that we deem closest to the current-stage prior $p_{t}(\boldsymbol{\theta})$ into the KL divergence. When the prior is a multivariate Gaussian distribution, this term can be computed in closed form as
$\mathrm{KL}\left(\mathcal{N}(\boldsymbol{\mu}_{1},\boldsymbol{\Sigma}_{1})\parallel\mathcal{N}(\boldsymbol{\mu}_{2},\boldsymbol{\Sigma}_{2})\right)=\frac{1}{2}\left(\text{tr}(\boldsymbol{\Sigma}_{2}^{-1}\boldsymbol{\Sigma}_{1})+(\boldsymbol{\mu}_{2}-\boldsymbol{\mu}_{1})^{\top}\boldsymbol{\Sigma}_{2}^{-1}(\boldsymbol{\mu}_{2}-\boldsymbol{\mu}_{1})-k+\ln\left(\frac{\det(\boldsymbol{\Sigma}_{2})}{\det(\boldsymbol{\Sigma}_{1})}\right)\right).$ (27)
where $\boldsymbol{\Sigma}=\text{diag}(\boldsymbol{\sigma^{2}})$, $k$ represents the dimensionality of the distributions, $\text{tr}(\cdot)$ denotes the trace of a matrix, and $\det(\cdot)$ stands for the determinant of a matrix. For the case that the prior is a mixture of Gaussian distributions, we can refer to the next section to get its upper bound.
Appendix C Mixture-of-Gaussian Prior
We refer to the lemma that was stated for the mixture of Gaussian in Singer & Warmuth (1998).
Lemma C.1.
The KL divergence between two mixture distributions $p=\sum_{i=1}^{k}\alpha_{i}p_{i}$ and $p^{\prime}=\sum_{i=1}^{k}\alpha_{i}p^{\prime}_{i}$ is upper-bounded by
Proof.
Using the log-sum inequality (Cover, 1999), we have
$\leq\int\sum_{i=1}^{k}\alpha_{i}p_{i}\log\cfrac{\alpha_{i}p_{i}}{\alpha_{i}p^{\prime}_{i}}$
$=\sum_{i=1}^{k}\alpha_{i}\left(\int p_{i}\log\cfrac{\alpha_{i}}{\alpha_{i}^{\prime}}+\int p_{i}\log\cfrac{p_{i}}{p^{\prime}_{i}}\right)$
So the lower bound (26) can be redefined as
$\mathcal{L}$ $=-\lambda\mathbb{E}_{\boldsymbol{\theta}\sim q(\boldsymbol{\theta})}H(\mathcal{U}_{t}|\boldsymbol{\theta})-\mathrm{KL}\left(q(\boldsymbol{\theta})\parallel p_{t}(\boldsymbol{\theta})\right)$ (30)
$\geq-\lambda\mathbb{E}_{\boldsymbol{\theta}\sim q(\boldsymbol{\theta})}H(\mathcal{U}_{t}|\boldsymbol{\theta})-\alpha\cdot\text{KL}\left(q||{p}_{1}\right)-(1-\alpha)\cdot\text{KL}\left(q||\bar{p}_{t}\right)\stackrel{{\scriptstyle\text{def}}}{{=}}\mathcal{L}^{\prime},$
Then, we have obtained a lower bound that can be optimized through closed-form calculations as the source prior distribution $q_{0}(\boldsymbol{\theta})$ and the teacher prior distribution $\bar{q}_{t}(\boldsymbol{\theta})$ are multivariate Gaussian distributions, which means we can also optimize $\mathcal{L}^{\prime}$ with Eq. (27).
Appendix D Recursive variational approximation process in VCoTTA
In this section, we construct a complete algorithmic workflow utilizing variational approximation in VCoTTA.
Before test time: First, we adopt a variational warm-up strategy to inject stochastic dynamics into the model before adaptation. Given the source dataset $\mathcal{D}_{0}$, we can use a variational approximation of $p(\boldsymbol{\theta}|\mathcal{D}_{0})$ as follows
$p(\boldsymbol{\theta}|\mathcal{D}_{0})=p_{1}(\boldsymbol{\theta})\approx q_{0}(\boldsymbol{\theta})=\arg\min_{q\in\mathbb{Q}}\mathrm{KL}\left[q(\boldsymbol{\theta})\parallel\frac{1}{Z_{0}}p(\boldsymbol{\theta})p(\mathcal{D}_{0}|\boldsymbol{\theta})\right],$ (31)
where we use the pretrained deterministic model $p_{0}(\boldsymbol{\theta})$ as the prior distribution.
When the domain shift: Then, at the beginning of the test time, we set the prior in task $t$ as $p_{t}(\boldsymbol{\theta})=\alpha\cdot p_{1}(\boldsymbol{\theta})+(1-\alpha)\cdot\bar{p}_{t}(\boldsymbol{\theta})$ and variational approximation, where $p_{1}(\boldsymbol{\theta})\approx q_{0}(\boldsymbol{\theta})$ and $\bar{p}_{t}(\boldsymbol{\theta})\approx\bar{q}_{t}(\boldsymbol{\theta})$. For $\bar{q}_{t}(\boldsymbol{\theta})$, which means the real-time posterior probability of the teacher model for the $t$-th test domain, is constantly updated by $q_{t}(\boldsymbol{\theta})$ via EMA (see Sec. 4.3.3) during the test phase. Note that we do not have $\bar{q}_{t}(\boldsymbol{\theta})$ for the first update in the $t$-th phase. In fact, we use $q_{t-1}(\boldsymbol{\theta})$ construct the prior. In this scenario, we use $p_{t}(\boldsymbol{\theta})\approx\alpha\cdot p_{1}(\boldsymbol{\theta})+(1-\alpha)\cdot q_{t-1}(\boldsymbol{\theta})$. This is the variational distribution that should be used to approximate the prior in the absence of a teacher model in the first step, as well as the approximation that should be used when not employing the MT architecture. That is, using the posterior approximation from the previous domain to approximate the prior of the next domain is somehow like Nguyen et al. (2018). Note that the process is not required to inform the model that the domain produces a shift.
During a test domain: With the approximation to $p_{t}(\boldsymbol{\theta})$ and analysis from Appendix A, we can get $q_{t}(\boldsymbol{\theta})$ for student model at the test domain $t$ as follows:
which means, we can recursively derive $p_{t+1}(\boldsymbol{\theta})$ and the following variational distributions, thereby achieving the goal of VCoTTA.

### v2

@@ VI-G1 Confidence margin in augmentations
VI-G2 Different number of augmentation
In our method, we also use augmentation to enhance the confidence. We then evaluate the the number of augmentation in Eq. (11). The results can be seen in Table VII, and shows that increasing the number of augmentations can enhance effectiveness, but this hyperparameter ceases to have a significant impact after reaching 32.
TABLE VII: Different number of augmentation.
@@ VI-G2 Different number of augmentation
VI-H Analysis of Variational Warm-up Strategy
We have discussed the Variational Warm-Up (VWU) strategy in Section IV-C1, and explain that the warm-up strategy is a common practice in TTA and CTTA. We further illustrate some attributes of the proposed variational warm-up strategy. In our method, the VWU strategy is used to turn an off-the-shelf model to a pretrained BNN. The advantage of this approach is that pretrained models are readily available (e.g., directly leveraging official models in PyTorch), while pretrained BNNs are challenging to obtain, especially for large-scale datasets. Moreover, training BNNs is more difficult. Therefore, constructing BNN pretrained models based on existing pretrained models is a feasible approach. Additionally, we find that such a warm-up strategy requires only a few epochs to achieve satisfactory results. To validate the characteristics of the proposed VWU strategy, we designed the following experiments.
VI-H1 Warm-up vs. Directly Pretraining BNN
First, we conducted experiments to compare the performance of obtaining pretrained BNN models using the warm-up approach versus directly training the source model with BNN. We pretrain the BNN also use VI as describing in Section IV-C1. The results can be seen in Fig. 4. As we can see, the results are at the same level, for example VI pretraining is with 13.2% error rate while the proposed VWU achieves 13.1% on CIFAR10C. However, if we direct turn a pretrained model to a BNN by adding random stochastic parameters, without warm-up strategy, the results drop to 17.1%. This shows that VWU is a feasible strategy to obtain a pretrained BNN.
Figure 4: Comparison of variational warm-up and directly BNN pretraining.
VI-H2 Number of Warm-up Epochs
In our implementation, we employ only a limited number of epochs for variational warm-up, say 5 epochs. This is due to the fact that the pretrained model already fits well, thus requiring minimal adjustments to the mean of BNN. Additionally, the standard deviation (std) is initialized to be small. Consequently, only a small number of iterations are necessary to update the BNN, and the step size is also kept small. Experimentation on the epoch number of variational warm-up reveals that keeping increasing epochs ( >5>5) will diminish performance, as shown in Fig. 5(a).
(a) Different warm-up epochs.
(b) Different warm-up data scale.
Figure 5: Comparisons on different warm-up settings of CIFAR10C.
VI-H3 Only Portion Usage of Source Dataset in Warm-up
The warm-up strategy is a common approach in TTA and CTTA tasks and it is regarded as a part of the pretraining stage. We also evaluate how we only use partial data for warm-up, and the results are shown in Fig. 5(b). The experimental results demonstrate that a moderate reduction in sample size still maintains certain effectiveness of the warm-up strategy. However, excessive reduction, such as reducing to $1/10$, leads to a certain decline in effectiveness. This is because the warm-up strategy aims to incorporate statistical information from the dataset into the model, and insufficient data may result in inaccurate performance.
VI-I Different Orders of Corruption
As we discuss in the major comparisons (see Sec VI-B), the performance may be affected by the corruption order. To provide a more comprehensive evaluation of the matter of the order, we conduct 10 different orders from Sec VI-B, and show the average performance of all compared methods. 10 independent random orders of corruption are all under the severity level of 5. The results are shown in Table VIII. We find that the order of corruption is minor on simple datasets such as CIFAR10C and CIFAR100C, but small std on difficult datasets such as ImageNetC. The proposed VCoTTA outperforms other methods on the average error of CIFAR10C and CIFAR100C under 10 different corruption orders, which shows the effectiveness of the prior calibration in CTTA. Moreover, VCoTTA has comparable results with PETAL on ImageNetC, but smaller std over 10 orders, which shows the robustness of the proposed method.
TABLE VIII: Comparisons over 10 orders (avg $\pm$ std).
Method CIFAR10C CIFAR100C ImageNetC
@@ VI-I Different Orders of Corruption
[…]
TABLE IX: Resetting to the source model when domain shifts.
Method Reset C1 C2 C3 C4 C5 C6 C7 C8 C9 C10 C11 C12 C13 C14 C15 avg
CoTTA 40.1 37.7 39.7 26.9 38.0 27.9 26.4 32.8 31.8 40.3 24.7 26.9 32.5 28.3 33.5 32.5
CoTTA $\surd$ 40.1 40.2 42.5 27.4 41.6 29.5 27.6 34.6 34.8 41.2 26.4 30.0 35.3 32.6 40.8 35.1
VCoTTA 35.3 32.8 38.9 23.8 34.6 25.5 23.2 27.5 26.7 30.4 22.1 23.0 28.1 24.2 30.4 28.4
VCoTTA $\surd$ 35.3 33.7 38.7 24.0 37.0 26.2 24.1 29.2 28.6 33.9 22.2 23.7 30.9 26.2 34.5 29.9
VI-J Corruption Loops
Figure 6: 10 loops under a same corruption order (CIFAR10C).
VI-K Experiment on different batch sizes
CTTA does operate in an online setting, where all testing data is used only once. We evaluate our methods with different batch sizes. For a testing dataset with a fixed number of samples, a larger batch size means fewer updates. The current focus of CTTA research primarily revolves around batch-mode online settings, with batch sizes typically set to 200 in our experiments like other SOTAs. Moreover, strict online learning settings where each data point is processed individually are under-researched. In fact, our method can be applied in scenarios with online learning or small batch sizes. We experiment with the batch size of 1 on CIFAR10C, and compare the results with some baseline methods. The comparison results are shown in Fig. 7. Note that the batch normalization (BN) layers are disabled when the batch size is 1. The results show that small batch sizes in CTTA result in worse performance. We believe this is because a small batch size amplifies the uncertainty in model training. We also find that TENT has high errors when the batch size is small but larger than 1. This is because TENT only updates the BN layers, which highly depends on the size of mini batch.
Figure 7: Error comparisons of different test batch sizes.
VI-L Can model reset mitigate the error accumulation issue?
In the experiments of CTTA, it is interesting to know if resetting to source simply can solve the drift issue and achieve better results than CTTA methods. We conduct this experiment on CoTTA and our method, the results can be seen in Table IX. We find that the resetting performance is slightly less effective than the continual setting on both CoTTA and our method. The results indicate that there may be shared knowledge between different tasks in the CTTA. Directly resetting to the source model would overlook this shared knowledge. Although the issue of error accumulation no longer needs to be considered, it becomes challenging to effectively mine some of the shared information.
VI-M Time and Memory Cost
We implement our method using a single RTX-4090 GPU card. We provide the memory and time cost on CIFAR10C in Table X. Our proposed VCoTTA method does not offer an advantage in terms of memory usage. This is because, in the BNN framework, additional standard deviations are required for implementing local reparameterization tricks. However, during the testing phase, this does not significantly impact the efficiency of the model. This is because, during testing, only the student model employs variational inference, which requires uncertainty parameters.
TABLE X: Time and memory cost comparisons.
Method Memory Time per corruption
CoTTA 10.3Gb 272s
PETAL 10.2Gb 261s
VCoTTA 11.1Gb 279s
VI-N Different learning rate
We also adjust the learning rate and select suitable learning rates for different datasets. As shown in the figure, through experimentation, we choose learning rates of 1e41e-4, 1e41e-4, and 1e21e-2 for the CIFAR10C, CIFAR100C, and ImagenetC datasets, respectively. Note that this result is only applicable to the experimental scenario. It is challenging to select the optimal learning rate in a real-world testing environment, as the future testing scenarios are unknown.
Figure 8: Learning rate analysis of the proposed method.
VII Conclusion and Limitation
In this paper, we proposed a variational Bayesian inference approach, termed VCoTTA, to estimate uncertainties in CTTA. At the pretrained stage, we first transformed an off-the-shelf pretrained deterministic model into a BNN using a variational warm-up strategy, thereby injecting uncertainty into the source model. At the test time, we implemented a mean-teacher update strategy, where the student model is updated via variational inference, while the teacher model is refined by the exponential moving average. Specifically, to update the student model, we proposed a novel approach that utilizes a mixture of priors from both the source and teacher models. Consequently, the ELBO can be formulated as the cross-entropy between the student and teacher models, combined with the KL divergence of the prior mixture. We demonstrated the effectiveness of the proposed method on three datasets, and the results show that the proposed method can mitigate the issue of unreliable prior within the CTTA framework.
The efficacy of the proposed method relies on injecting uncertainty into the model during the pre-training phase, which may be unavailable in scenarios where pretraining is already completed, and original data is inaccessible. Additionally, constructing and training BNN models are inherently more complex, highlighting the importance of enhancing computational efficiency. The Gaussian mixture method relies on multiple data augmentations, which also incurs computational costs. Future endeavors could explore more efficient approaches for Gaussian mixtures.